OpenStudy (anonymous):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

(.7)x = 500 The .7 represents the sales price of each flower. The x represents how many flowers are required to make a profit of whatever is on the right side of the equation. The 500 represents the desired profit. \[.7x = 500 \\ x = 714.285714\] If you want me to explain how I got the .7, just let me know.

OpenStudy (anonymous):

thank you for answering! and yes could you please explain

OpenStudy (anonymous):

@ColossalMagnitude

OpenStudy (anonymous):

Okay. Part of the sales price is the purchase price, which is $0.50. Another part of the sales price is the 40% of the purchase price, which is 40% * $0.50, which is .4 * $0.50, which is $0.20. $0.50 + $0.20 is $0.70, which is how I got .7. Hope this helps!

OpenStudy (anonymous):

what do you mean when you say whatever is on the right side of the equation?

OpenStudy (anonymous):

@ColossalMagnitude

OpenStudy (anonymous):

The right side of the equation represents the desired profit, so I was basically just saying that x represents how many flowers are required to make the desired profit.

OpenStudy (anonymous):

instead of saying what is on the right side is their something else I could put for my answer?

OpenStudy (anonymous):

Instead of saying "The x represents how many flowers are required to make a profit of whatever is on the right side of the equation.", just say "The x represents how many flowers are required to make the desired profit."

OpenStudy (anonymous):

@ColossalMagnitude do ya think you could help me with another?

OpenStudy (anonymous):

Yeah, sure.

OpenStudy (anonymous):

Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10 and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (anonymous):

@ColossalMagnitude for the first question the answer is 7?

OpenStudy (anonymous):

Um, can you define what you mean be first question?

OpenStudy (anonymous):

What is the answer for the first question?

OpenStudy (anonymous):

The one about the flower fundraising?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ColossalMagnitude

OpenStudy (anonymous):

Sorry, my Internet Connection is complete crap *excuse the "profanity"*. The answer is 500 / .7, or 714.285714.

OpenStudy (anonymous):

thank you :3

OpenStudy (anonymous):

could you help with this question please? A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

hold on...I think #1 is wrong...let me check it..

OpenStudy (anonymous):

alright I appreciate it thank you

OpenStudy (anonymous):

|x − 0.525| = 0.01 We get the following equations from this absolute-value: x - 0.525 = 0.01 x = 0.535 x - 0.525 = -0.01 x = .0515 The maximum diameter is 0.535. The minimum diameter is .0515.

OpenStudy (anonymous):

No, scratch that, the minimum diameter is 0.515.

OpenStudy (texaschic101):

It costs them .50 to purchase the flowers It costs 40% of 50, which is .70 to sell. This each flower makes them a profit of .20 .20x = 500 x = 500/.20 x = 2500 check... it will cost them .50 (2500) = 1250 they sell them for .7(2500) = 1750 profit : 1750 - 1250 = 500 (correct) They need to sell 2500 flowers to make a profit of $500

OpenStudy (anonymous):

thank you could you answer this last question? A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

Um, I think you misunderstood that question @texaschic101 . The sales price of each flower is the purchase price, which is $0.50, PLUS the 40% sales tax, which is $0.20. So, in total, it would be $0.70. This means the equation would actually be: .7x = 500 x = 500/.7 x = 714.285714. They need to sell 715 flowers to make a profit of $500. Also, at @MyChem, I already answered that assembly line question.

OpenStudy (texaschic101):

they are selling the flowers for .70...you are right about that, but you are forgetting to subtract what they paid for the flowers, which is .50 per flower. Thus, giving them .20 profit on each flower sold. Profit = selling price - purchase price. So if they purchased 2500 flowers, it would cost them 1250. And by selling them at .70 each, they make 1750. Thus, making them a profit of $500.

OpenStudy (anonymous):

Oh, I am so sorry. You are completely right, how could I have forgotten to have done that. *facepalm*

OpenStudy (texaschic101):

lol....If I had $1 for every mistake I made, I would be rich :)

OpenStudy (anonymous):

do yall think you could help me with my world history?

OpenStudy (texaschic101):

I can try, but you would probably get more help in the history section.....although there is not always that many people in the history section

OpenStudy (anonymous):

I took world history a while back; I don't think I could remember that stuff, sorry.

OpenStudy (texaschic101):

I suggest posting it in the history section or at least post a new question because this post is beginning to look like a book...lol